OpenStudy (anonymous):

prove the identity by use of the following fundamental identities. csc^2ytan^2y=tan^2y+1

OpenStudy (anonymous):

LHS =1/ sin^y . sin^2 y/cos^2y =1/cos^2y =(sin^2y+cos^2y)/cos^y =tan^2y+1 =RHS

OpenStudy (anonymous):

csc^2 y tan^2 y - tan^2 y = 1 tan^2 y (csc^2 y - 1) = 1 csc^2 y - 1 = cot ^2 y which is a pythagorean identity :)

OpenStudy (anonymous):

so the answer here is correct http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet_Reduced.pdf

OpenStudy (anonymous):

what is rhs? haha

OpenStudy (anonymous):

Right hand side lol !!

OpenStudy (anonymous):

can your solution will be the answer?

OpenStudy (anonymous):

yes !

OpenStudy (anonymous):

thx!

OpenStudy (anonymous):

another 1 i post it hehe

OpenStudy (anonymous):

(cotx+1)^2=csc^2x+2cotx

OpenStudy (anonymous):

LHS cot^2x+2cotx+1 since 1+cot^2x=csc^2x csc^2x+2cotx RHS

OpenStudy (anonymous):

what is rhs?

OpenStudy (anonymous):

LHS - Left Hand Side RHS - Right Hand Side

OpenStudy (anonymous):

since LHS=RHS, instead of writing the term i just put LHS

OpenStudy (anonymous):

ic ic thx

OpenStudy (anonymous):

so what is the purpose of rhs in my question?

OpenStudy (anonymous):

u have given both LHS term RHS term, u want to prove how LHS term equals RHS term.is n't it?

OpenStudy (anonymous):

ahhh i get it now

OpenStudy (anonymous):

but if i write it on my h.w.will i put it or it is ok that rhs/lhs is not there?

OpenStudy (anonymous):

In some books, the proof of any such things is given in LHS/RHS form. But for ur case better ask ur teacher.

OpenStudy (anonymous):

last 1 sir \[{1-\cos ^{2}\beta}\over {\cos \beta} \] =sin B tan B

OpenStudy (anonymous):

LHS =((sin^2B+cos^2B)-cos^2B)/cosB =sin^2B/cosB =sinB.sinB/cosB =sinB tanB RHS

OpenStudy (anonymous):

thx sir your the best last pls (cscx-cotx) (cscx+cotx)=1

OpenStudy (anonymous):

LHS =csc^2x-cot^2x =1 RHS

OpenStudy (anonymous):

thx again gowthaman

OpenStudy (anonymous):

(sin^2+cos^2)=1+2sin2cos2

OpenStudy (anonymous):

something wrong

OpenStudy (anonymous):

yes theres is w8

OpenStudy (anonymous):

(sin^2+cos^2)^2=1+2sin2cos2

OpenStudy (anonymous):

still something is wrong it should be (sinx + cosx)^2 = 1 + 2sinx cosx or = 1+sin2x

OpenStudy (anonymous):

i thought its 2 sory again >>> here (sinx + cosx)^2 = 1 + 2sinx cosx

OpenStudy (anonymous):

expand it like (a+b)^2=a^2+b^2+2ab

OpenStudy (anonymous):

in the last part is that 2sinxcosx?

OpenStudy (anonymous):

2AB=2SINXCOSX?

OpenStudy (anonymous):

S

OpenStudy (anonymous):

s?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

sory lag thx gowthaman

OpenStudy (anonymous):

my sister have this homework do you know some of this? obtain the positive solution less than 360 degree 1. sec^2 x-4=0 2. cot^2x=3 3. 2cos^2x=cosx 4. cot^2x=cotx 5. 4sin^2x-3=0

OpenStudy (anonymous):

its like some trigonometry angle or something quadrant

OpenStudy (anonymous):

1. sec^2x-4=0 sec^2x=4 secx=+/- 2 or cosx = +/- 1/2 I think now u can solve for x